OpenStudy (anonymous):

Solve 5x - 6y = -38 3x + 4y = 0

OpenStudy (anonymous):

x=-4 , y=3

OpenStudy (shamil98):

please show your work, and don't just give out the answer..

OpenStudy (anonymous):

Sorry. one sec. I'll type it up.

OpenStudy (shamil98):

2(5x - 6y = -38) 3(3x + 4y = 0) you could solve this through elimination. multiply both equations to get rid of y 10x - 12y = -76 9x + 12y = 0 add the equations ------------ 19x = -76 x= 4 Plug in x in the equation to solve for y 3(4) + 4y = 0 12 + 4y = 0 4y = -12 y= -3

OpenStudy (anonymous):

Thank You ^_^

OpenStudy (shamil98):

Yeah, no problem.

OpenStudy (anonymous):

thats not right.

OpenStudy (shamil98):

oh oops, i meant -4,3. -76/19 = -4. my bad.